A full joint distribution for the Toothache, Cavity, and Catch is given in the table below:

What is the probability of Cavity, given evidence of Toothache?
Pick an option to check yourself, or open the answer below.
Show answer & explanation
Correct answer: D. <0.6,0.4>
Explanation
P(toothache) = 0.108 + 0.012 + 0.016 + 0.064 = 0.2; P(cavity ∧ toothache) = 0.12. So P(Cavity | toothache) = ⟨0.12/0.2, 0.08/0.2⟩ = ⟨0.6, 0.4⟩.
More Artificial Intelligence (AI) questions
- In Artificial Intelligence (AI), an environment is uncertain if it is ________.
- In Artificial Intelligence (AI), a simple reflex agent selects actions on the basis of_________.
- In heuristic search algorithms in Artificial Intelligence (AI), if a collection of admissible h … … . h heuristics 1 m is available for a p…
- Consider the following sentences about A* search: (a) A* expands all nodes with f(n) < C*. (b) A* expands no nodes with f(n) ≥ C*. (c) Prun…
- Consider a vocabulary with only four propositions A, B, C and D. How many models are there for the sentence B ∨ C?
- Consider the following statements: (a) False ⊨ True (b) If α ⊨ (β ∧ γ) then α ⊨ β and α ⊨ γ. Which of the following is correct?